OpenStudy (anonymous):

What is the answer to 2(3x+5)-21=4x-7?

OpenStudy (anonymous):

x = 2

OpenStudy (dinamix):

x=2

OpenStudy (zzr0ck3r):

\[2(3x+5)-21=4x-7\\6x+10-21=4x-7\\2x-11=-7\\2x=4\\x=2\]

OpenStudy (anonymous):

Thank you so much! I was having trouble because I haven't solved an equation like that in a while ! :)

OpenStudy (zzr0ck3r):

@robtobey @dinamix how does that help learn?

OpenStudy (zzr0ck3r):

Do you understand the steps?

OpenStudy (anonymous):

Yes thank you

OpenStudy (anonymous):

@zzr0ck3r The problem poster requested, "What is the answer"

OpenStudy (zzr0ck3r):

lol